OpenStudy (anonymous):

A parallelogram has an area of 20.3 square centimeters and a base of 2.5 centimeters. Write an equation that relates to the base and height. Solve the equation to determine the length of the height.

OpenStudy (anonymous):

a=b*h, therefore h=a/b or b=a/h, 20.3=2.5*h.

OpenStudy (anonymous):

I understand know. Thanks!